A telemarketer with B2B Communications has a 20 percent historical probability of making a sale during a shift where 100 calls were made. Assign a random variable x, where the value of x is equal to the number of sales made during the shift. Describe the probability distribution of x, and calculate the probability of the employee making exactly 28 sales during a 100-call shift. What is the probability that the employee will make less than 28 sales during the 100-call shift? What is the probability that the employee will make 28 or more sales during the 100-call shift?

Answer by stanbon(75887)

A telemarketer with B2B Communications has a 20 percent historical probability of making a sale during a shift where 100 calls were made.
Assign a random variable x, where the value of x is equal to the number of sales made during the shift.
------------------------------------
Describe the probability distribution of x, and calculate the probability of the employee making exactly 28 sales during a 100-call shift.
mean = np = 100*0.20 = 20
std = sqrt(npq) = sqrt(20*0.8) = 4
------------------------------------------
What is the probability that the employee will make less than 28 sales during the 100-call shift?
P(0<= x <=27) = binomcdf(100,0.2,27) = 0.9658
------
What is the probability that the employee will make 28 or more sales during the 100-call shift?
P(28<= x <=100) = 1-0.9658 = 0.0342
=========================================
Cheers,
Stan H.
=============